A student sits on a rotating stool holding two
2 kg objects. When his arms are extended
horizontally, the objects are 1 m from the axis
of rotation, and he rotates with angular speed
of 0.61 rad/sec. The moment of inertia of the
student plus the stool is 6 kg m2
and is assumed to be constant. The student then pulls
the objects horizontally to a radius 0.27 m
from the rotation axis Calculate the final angular speed of the
student.
Answer in units of rad/s. Don't round answer.
1
Expert's answer
2020-01-20T05:21:21-0500
"I_i=6+2(2)(1)^2=10\ kg\cdot m^2"

"I_f=6+2(2)(0.27)^2=6.29\ kg\cdot m^2"

"\omega_f=\omega_i\frac{I_i}{I_f}=0.61\frac{10}{6.29}=0.97\frac{rad}{s}"